How can you graph 6.5x+10y=-20 and -6x+3y=18
Answer by mananth(16946)

6.5x+10y=-20
6.5 x + 10 y = -20
y= - 13/20 x -2
Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 20 -20
y= -2 -15 11
(x,y) (0,-2),(20,-15),(-20,11 )



-6x+3y=18
-6 x + 3 y = 18
y= 2 x + 6

Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 1 2
y= 6 8 10
(x,y) (0,6),(1,8),(2,10)



The solution is (-3,0)